OpenStudy (anonymous):

Tell whether the function y=6(x)^5 is exponential. Explain your answer.

OpenStudy (anonymous):

exp means x^a , where x is a variable > 0 and a is a constant. Hence it an exponential.

OpenStudy (anonymous):

Exponential is a function of the form y = a^x. The form y=cx^a corresponds to a Power function.